OpenStudy (anonymous):

Show that the equation of the line passing through the points (3,2) and (-4,16) has the equation y=-2x+8

OpenStudy (the_fizicx99):

I think they're asking you if after you solve it, if it gives you y = -2x + 8 Well first find the slope, M = y2-y1/x2-x1 M = 16-2/-4-2 M = 14/-6 Then plug into y-y1=M(x-x1) y-2 = -14/6(x - 3) y - 2 = -14/6x + 7 +2 +2 y = -14/6x + 9

OpenStudy (anonymous):

Well first find the slope byusing y2-y1/x2-x1 you can use the points either way. I'll just use 16-2/-4-7 = 14/-7 ; m = -7 SO now that we have the slope you can just put it into point slope form and simplify. In point slope form, you can substitute either values for y1 and x1. Point slope form = y - y1= m(x-x1) Use the first points (3,2) Now substitute them in for y1 and x1 and substitute in your slope. y - 2 = -2(x-3) Now just simplify. y - 2 = -2x + 6 y - 2 + 2 = -2x +6 + 2 y= -2x + 8 x

OpenStudy (anonymous):

Thank you

OpenStudy (the_fizicx99):

@sarara your answer is incorrect.

OpenStudy (anonymous):

then can you help tHe_fiZ

OpenStudy (anonymous):

oh why?

OpenStudy (the_fizicx99):

I think I also have a typo, M = 16-2/-4-3 = 14/-7 = -2 y-2 = -2(x-3) y -2 = -2x + 6 +2 +2 y = -2x + 8

OpenStudy (the_fizicx99):

You also had the same typo as I did, you had m = -7

OpenStudy (anonymous):

I meant to put m = -2. But I'm pretty sure it's right cause I tried both points.

OpenStudy (anonymous):

Thank you both for your help. I just needed the explanation of how to do it and you both helped. Thanks

OpenStudy (the_fizicx99):

both points?

OpenStudy (anonymous):

yeah i understand how to do it now Thanks